Filer Manufacturing has 7.3 million shares of common stock outstanding. The current share price is $43, and the book value per share is $3. The company also has two bond issues outstanding. The first bond issue has a face value of $68 million and a coupon rate of 6 percent and sells for 109.3 percent of par. The second issue has a face value of $58 million and a coupon rate of 6.5 percent and sells for 106.9 percent of par. The first issue matures in 7 years, the second in 28 years.

Suppose the company’s stock has a beta of 1.4. The risk-free rate is 2.1 percent, and the market risk premium is 6 percent. Assume that the overall cost of debt is the weighted average implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 34 percent. What is the company’s WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  

  WACC %
Answer #1
Solution:
WACC =8.35%
Working Notes:
Common stock = 7.3 million shares
Bond 1 = Face value of $68 million
Bond 2 = Face value of $58 million
Current share price = $43
Bond 1 is selling at 109.3% of par
Bond 2 is selling at 106.9% of par
Total Market value of common stock (E) = No. of Common stock shares x Market price per share
Total Market value of common stock (E) = 7.3 million x $43
Total Market value of common stock (E) = $313.9 million
Total Market value of Bond 1 = Total Face value of bond x % of par at which bond is selling in market
Total Market value of Bond 1 = $68 million x 109.3%
Total Market value of Bond 1 = $74.324 million
Total Market value of Bond 2 = Total Face value of bond x % of par at which bond is selling in market
Total Market value of Bond 2 = $58 million x 106.9%
Total Market value of Bond 2 = $62.002 million
Total debt market value(D) =Bond 1 +Bond 2 market value
Total debt market value(D) =$74.324 million   + $62.002 million
Total debt market value(D) =$136.326 million
the firm’s market value company capital structure (V) = E + D = $313.9 million + $136.326 million
the firm’s market value company capital structure (V) = E + D = $450.226 million
Debt (Bond) weight in capital structure = D/V = Mkt. Value of Bond / Total Mkt. Value of Company
Debt (Bond) weight in capital structure = $136.326 million /$450.226 million
Debt (Bond) weight in capital structure = 0.302794596
Common stock weight in capital structure = E/V = Mkt. Value of common stock / Total Mkt. Value of Company
Common stock weight in capital structure = E/V = $313.9 million /$450.226 million
Common stock weight in capital structure = E/V =0.697205404
Cost of Equity (Ke)
Using CAPM
r= rf + (rm-rf) x B
where
r =Cost of Equity (Ke)= ???
rf= risk free rate = 2.1%
(rm -rf)= Market risk premium = 6%
B= beta=1.4
r= rf + (rm-rf) x B
r= 2.1% + 6% x 1.4
r= 2.1% + 8.4%
r= Cost of Equity (Ke) =10.5%
cost of debt pre tax (kd)
Total debt market value(D) =$136.326 million
Total debt market value(D) =Bond 1 +Bond 2 market value
Total debt market value(D) =$74.324 million   + $62.002 million
weight of bond 1 in total debt d1= mkt. Value of Bond 1/total debt = $74.324 / $136.326
weight of bond 1 in total debt d1=0.54519314
weight of bond 2 in total debt d2= mkt. Value of Bond 2/total debt = $62.002 / $136.326
weight of bond 2 in total debt d2=0.45480686
Computation of YTM of Bond 1
As the bond is paying coupon semi annually , its Ytm can be calculated by Excel or financial calculator
First we get the semi annual YTM
No. of period = years to maturity x no. of coupon in a year = 7 x 2 =nper = N = 14
Face value of bond = FV= $1,000
Price of the bond = PV = -$1093         [1000 x 109.3% = $1093]
Semi-annual Coupon amount = PMT = coupon rate x face value/2 = 6% x $1,000 /2= $30
For calculation YTM by excel
type above data in below format
=RATE(N,pmt,PV,FV)
=RATE(14,30,-1093,1000)
2.2198646%
=2.2198646%
The YTM calculated is semi annual
YTM annual = Semi annual YTM x 2
YTM annual = 2.2198646% x 2
YTM annual bond 1 = 4.439729139 %
Computation of YTM of Bond 2
As the bond is paying coupon semi annually , its Ytm can be calculated by Excel or financial calculator
First we get the semi annual YTM
No. of period = years to maturity x no. of coupon in a year = 28 x 2 =nper = N = 56
Face value of bond = FV= $1,000
Price of the bond = PV = -$1069         [1000 x 106.9% = $1069]
Semi-annual Coupon amount = PMT = coupon rate x face value/2 = 6.5% x $1,000 /2= $32.50
For calculation YTM by excel
type above data in below format
=RATE(N,pmt,PV,FV)
=RATE(56,32.5,-1069,1000)
2.99441020048196%
=2.9944102%
The YTM calculated is semi annual
YTM annual = Semi annual YTM x 2
YTM annual = 2.9944102% x 2
YTM annual bond 2 = 5.9888204%
Cost of debt (Kd) = (weight of bond 1 x YTM of bond1) + (weight of bond 2 x YTM of bond2)
weight of bond 1 in total debt d1=0.54519314
weight of bond 2 in total debt d2=0.45480686
YTM annual bond 1 = 4.439729139 %
YTM of bond2 = 5.9888204%
Cost of debt (Kd) = (weight of bond 1 x YTM of bond1) + (weight of bond 2 x YTM of bond2)
Cost of debt (Kd) = (0.54519314 x 4.439729139 %) + (0.45480686 x 5.9888204%)
Cost of debt (Kd) = 0.051442665
Cost of debt (Kd) = 5.14426647%
After Tax cost of debt (kd)= Kd x (1 - tax rate) = 5.14426647% x (1-0.34)
After Tax cost of debt (kd)=3.39521587%
WACC
WACC = (E/V x Ke) + (D/V x After tax Kd)
= (0.697205404 x 10.50% +0.302794596 x 3.39521587%)
0.083487098
=0.0835
=8.35%
WACC =8.35%
Where
Debt (Bond) weight in capital structure = 0.302794596
Common stock weight in capital structure = E/V =0.697205404
r= Cost of Equity (Ke) =10.5%
After Tax cost of debt (kd)=3.39521587%
Please feel free to ask if anything about above solution in comment section of the question.